How many grams of fluorine gas are in 3.28x10^5 L at a pressure of 5.79 kPa and a temperature of 46° C
1
Expert's answer
2020-03-23T10:13:54-0400

"PV=nRT"

"n=PV\/RT"

"V=3.28*10^5 L = 3.28 * 10^2 m^3"

"T= 319 K"

"P=5.79*10^3 Pa"

"n=5.79*10^3*3.28*10^2\/(319*8.314)=" "716 mole"

"m=n*M=716mole*38g\/mole=27208 g"

The answer is 27208g